### v1

# Detecting Functional Bugs in Smart Contracts through LLM-Powered and Bug-Oriented Composite Analysis

Abstract.
Smart contracts are fundamental pillars of the blockchain, playing a crucial role in facilitating various business transactions. However, these smart contracts are vulnerable to exploitable bugs that can lead to substantial monetary losses. A recent study reveals that over $80\%$ of these exploitable bugs, which are primarily functional bugs, can evade the detection of current tools. Automatically identifying functional bugs in smart contracts presents challenges from multiple perspectives. The primary issue is the significant gap between understanding the high-level logic of the business model and checking the low-level implementations in smart contracts. Furthermore, identifying deeply rooted functional bugs in smart contracts requires the automated generation of effective detection oracles based on various bug features.
To address these challenges, we design and implement PromFuzz, an automated and scalable system to detect functional bugs, in smart contracts. In PromFuzz, we first propose a novel Large Language Model (LLM)-driven analysis framework, which leverages a dual-agent prompt engineering strategy to pinpoint potentially vulnerable functions for further scrutiny. We then implement a dual-stage coupling approach, which focuses on generating invariant checkers that leverage logic information extracted from potentially vulnerable functions. Finally, we design a bug-oriented fuzzing engine, which maps the logical information from the high-level business model to the low-level smart contract implementations, and performs the bug-oriented fuzzing on targeted functions. We evaluate PromFuzz from $4$ perspectives on $6$ ground-truth datasets and compare it with multiple state-of-the-art methods. The results show that PromFuzz achieves $86.96\%$ recall and $93.02\%$ F1-score in detecting functional bugs, marking at least a $50\%$ improvement in both metrics over state-of-the-art methods. Moreover, we perform an in-depth analysis on $6$ real-world DeFi projects and detect $30$ zero-day bugs. Our further case studies, the risky first deposit bug and the AMM price oracle manipulation bug on real-world DeFi projects, demonstrate the serious risks of the exploitable functional bugs in smart contracts. Up to now, $24$ zero-day bugs have been assigned CVE IDs. Our discoveries have safeguarded assets totaling \$$\$18.2$ billion from potential monetary losses.
 † Binbin Zhao and Xingshuang Lin are the co-first authors. * Shouling Ji is the corresponding author.
1. Introduction
Blockchains have increasingly become a crucial component of the contemporary economic landscape. As of the writing of this paper, the total market capitalization of global cryptocurrencies has reached \$$\$3.5$ trillion (for, 2024). Smart contracts, which are fundamental pillars of the blockchain, facilitate the development of decentralized financial (DeFi) and various business transactions. Nevertheless, smart contracts are prone to exploitable bugs that often result in substantial financial losses. According to a report by CertiK, a leading Web3 security firm, over \$$\$1.8$ billion was lost due to $751$ security incidents in 2023 (cer, 2023).
Currently, there are many works have focused on detecting exploitable bugs in smart contracts, which can be categorized into four main types based on their methodologies: static analysis (Feist et al., 2019; Wang et al., 2019b), fuzzing (Choi et al., 2021; Shou et al., 2023), verification (So et al., 2020; Tan et al., 2022), and symbolic execution (myt, 2024; So et al., 2021). Despite these efforts, a recent study (Zhang et al., 2023) reveals that over 80% of exploitable bugs remain undetected by current tools, which are primarily functional bugs. This limitation mainly stems from the tools’ reliance on simple or hard-coded oracles to identify exploitable bugs. Nevertheless, detecting functional bugs involves comprehending the high-level business logic before scrutinizing the low-level implementations in smart contracts. Therefore, there is a pressing need for a practical system that can automatically detect functional bugs in smart contracts.
1.1. Challenges
To detect functional bugs in smart contracts automatically, we have the following key challenges.
Challenge I: Unraveling business logic in smart contracts. Detecting functional bugs in smart contracts requires a deep understanding of domain-specific properties intertwined with the contract’s business logic. The first challenge lies in automating the accurate extraction of this high-level business logic, a critical step for enabling functional bug detection. Designing an effective method to achieve this goal is difficult since business logic is often complex and intricately embedded within the low-level code implementations of smart contracts, posing significant challenges even for seasoned human auditors in distinguishing between business and code logic.
Challenge II: Bug checker generation. The business logic extracted from smart contracts is highly abstract and cannot be directly applied to functional bug detection. Thus, the second challenge involves deconstructing this business logic and creating practical bug checkers that can effectively utilize the extracted logic information. Designing bug checkers is non-trivial due to the varying fundamental characteristics exhibited by different types of functional bugs. Moreover, within diverse contract business contexts, the same type of functional bug may manifest different attributes, complicating the abstraction of functional bug features and the formulation of effective detection rules.
Challenge III: Functional bug detection. In practice, the effective deployment of bug checkers necessitates the support of sophisticated analysis tools for comprehensive bug detection. Currently, there is a gap in the availability of practical methods that can efficiently leverage these checkers. Therefore, the third challenge is to design a practical analysis method to efficiently utilize bug checkers for functional bug detection. Existing static analysis methods, while useful, struggle with false negatives as they cannot fully replicate the contract behavior in real-world environments, missing deep-level runtime-specific issues. On the other hand, traditional dynamic analysis methods explore programs randomly without a specific target, often relying on program crashes to identify bugs. However, functional bugs typically do not cause crashes and are unlikely to be triggered by random exploration alone.
1.2. Methodology
In this paper, we aim to address these challenges to detect functional bugs in smart contracts. To this end, we propose PromFuzz, an automated and practical system to conduct functional bug detection on smart contracts. Our design philosophy is as follows.
First, to solve the Challenge I, we start by designing a novel Large Language Model (LLM)-driven analysis framework. This framework is supported by a dual-agent prompt engineering strategy, featuring the Auditor Agent and the Attacker Agent. These specialized agents empower LLM to analyze smart contracts from two distinct perspectives: a meticulous auditor’s viewpoint and a malicious attacker’s perspective. We further enhance LLM’s capabilities by developing a results fusion algorithm, combining insights from both the Attacker Agent and the Auditor Agent. This integration aids in pinpointing potentially vulnerable functions for detailed examination, significantly enhancing the robustness of LLM’s output. Second, to address the Challenge II, we propose a dual-stage coupling approach, which focuses on generating invariant checkers that leverage logic information extracted from potentially vulnerable functions. In the initial stage, we implement a hierarchical matching approach that utilizes LLM’s capabilities to identify critical variables and principal statements relevant to various bug features. Subsequently, in the second stage, we implement a template-based checker generation method. This method involves designing 6 invariant checker templates capable of processing critical variables and principal statements to produce final invariant checkers. This dual-stage approach mitigates the randomness and hallucination inherent in LLM by breaking down complex tasks into simpler components, ensuring more reliable invariant checker generation. Third, to solve the Challenge III, we implement a bug-oriented analysis engine. The main idea behind our design is to strategically insert invariant checkers into specific potentially vulnerable functions, guiding the engine to explore these critical areas rather than searching aimlessly. By integrating invariant checkers into our analysis engine, we successfully detect functional bugs in smart contracts with high recall and F1-score.
1.3. Contributions
We summarize our main contributions as follows.
$\bullet$ We propose PromFuzz, an automated and practical system to detect exploitable functional bugs in smart contracts, which fills the gap between understanding the high-level business logic and scrutinizing the low-level implementations in smart contracts. We employ a novel dual-agent prompt engineering strategy that enables LLM to effectively extract hidden business logic from smart contracts. Additionally, our bug-oriented fuzzing engine successfully maps the logical information from the high-level business model to the low-level smart contract implementations.
$\bullet$ To the best of our knowledge, we build the first ground-truth dataset for verified functional bugs in smart contracts. This dataset includes 23 bugs from various DeFi projects. To facilitate future blockchain security research, we open-source both PromFuzz and the dataset at https://anonymous.4open.science/r/promfuzz-550D.
$\bullet$ We have implemented and evaluated PromFuzz across $4$ dimensions using $6$ ground-truth datasets. The results show that PromFuzz achieves $86.96\%$ recall and $93.02\%$ F1-score in detecting functional bugs. PromFuzz demonstrates its superiority in at least a $50\%$ improvement in both metrics over state-of-the-art methods.
$\bullet$ Our extensive analysis of $6$ real-world DeFi projects demonstrates PromFuzz’s great performance in detecting real-world functional bugs. Up to now, PromFuzz has identified $30$ zero-day bugs in $6$ real-world DeFi projects, with $24$ of these bugs have been assigned CVE IDs. Our discoveries have safeguarded assets totaling \$$\$18.2$ billion from potential monetary losses.
2. Background
In this section, we offer a succinct overview of the key concepts and techniques utilized throughout the paper, along with a motivating example.
2.1. Functional Bugs in Smart Contract
In this paper, we primarily focus on four major categories of functional bugs, including a total of ten subcategories of functional bugs, as shown in Table 1. These classifications cover the majority of functional bugs and are derived from prior research (Sun et al., 2024; Zhang et al., 2023).
Price Oracle Manipulation. In the blockchain ecosystem, smart contracts rely on price oracles to access real-world data like cryptocurrency prices. However, these oracles are susceptible to manipulation attacks, where attackers influence price data to impact transaction prices and gain illegal profits. We mainly focus on two specific price oracle manipulation bugs, Automated Market Maker (AMM) price oracle manipulation and non-AMM price oracle manipulation.
Unauthorized Behavior. Smart contracts often involve frequent token transfers, which, although a routine operation, are susceptible to potential attacks. These attacks stem from unauthorized behaviors—actions within a contract that deviate from its programmed rules or intended operational logic. Such bugs can result in loss of funds, corruption of the intended contract state, and other unintended effects that compromise the integrity and security of blockchain transactions. We specifically focus on two types of unauthorized behavior bugs: approval not clear and unauthorized transfer.
Insecure Calculating Logic. Smart contracts often exhibit vulnerabilities in their computational logic, which can lead to significant disruptions in their intended functions. A common issue arises when smart contracts do not accurately handle calculations related to share distributions, particularly during initial deposits. This can grant an unfair advantage to the first depositor, resulting in a disproportionate allocation of shares compared to subsequent participants. We focus on three types of insecure calculation logic bugs: wrong checkpoint order, wrong interest rate order, and risky first deposit.
Incorrect Control Mechanism. Smart contracts can suffer from critical flaws if they do not properly manage execution flows or restrict access to certain functionalities. Such mismanagement can trigger unintended behaviors or create vulnerabilities that attackers can exploit. This paper delves into three specific types of incorrect control mechanism bugs: improper handling of the deposit fee, wrong implementation of amount lock, and vote manipulation.
Table 1. Categories of functional bugs.
Bug Category Bug Subcategory
Price Oracle Manipulation AMM Price Oracle Manipulation
Non-AMM Price Oracle Manipulation
Unauthorized Behavior Approval Not Clear
Unauthorized Transfer
Insecure Calculating Logic Wrong Checkpoint Order
Wrong Interest Rate Order
Risky First Deposit
Incorrect Control Mechanism Improper Handling of the Deposit Fee
Wrong Implementation of Amount Lock
Vote Manipulation
2.2. Large Language Models
Large Language Models (LLMs) have demonstrated significant advancements across a variety of tasks, including code generation (Zheng et al., 2023), static analysis (Li et al., 2023), and program repair (Wei et al., 2023). Currently, a set of prominent LLMs have been developed, such as the Generative Pre-trained Transformer (GPT) by OpenAI, LLaMA (Touvron et al., 2023) by Meta, and Gemini (Anil et al., 2023) by Google. These models possess extensive world knowledge, strong problem-solving capabilities, sophisticated reasoning skills, and a robust capacity for instruction following.
LLMs exhibit significant emergent abilities (Wei et al., 2022a) that set them apart from traditional pre-trained models. They have three notable emergent abilities: in-context learning (Brown et al., 2020), instruction following (Ouyang et al., 2022), and step-by-step reasoning (Prystawski et al., 2023). For instance, small language models often struggle with complex tasks that require multi-step reasoning, such as analyzing the business logic concealed within smart contracts. In contrast, LLMs can adeptly manage these challenges using their emergent abilities. They utilize techniques like Chain-of-Thought (CoT) (Wei et al., 2022b) prompting to improve inference, thereby enabling them to address complex challenges effectively. In this paper, we utilize GPT-4-turbo as the foundational model in PromFuzz to analyze the hidden business logic in smart contracts. We focus on designing novel prompt strategies to trigger its emergent capabilities and implement practical measures to enhance its output robustness.
2.3. Motivating Example
We provide a functional bug discovered on Code4rena (Cod, 2024) as a motivating example, as shown in Figure 1. In the SimplePool contract, the transferFrom function facilitates the transfer of a specified amount of shares from the sender to the recipient using the balanceToShares function, which converts a specified token amount into shares based on the pricePerShare. However, a critical issue arises post-transfer: the approve operation erroneously subtracts the amountInShare instead of the amount. This mistake allows the approved spender potentially to transfer more tokens than intended, with the approved amount increasing in correlation with the pricePerShare.
A series of works, such as GPTScan (Sun et al., 2024) and SMARTINV (Wang et al., 2024), have explored the use of LLMs to detect functional bugs in smart contracts. GPTScan employs GPT-based matching analysis to initially identify potential functional bugs, which are then confirmed through static analysis. SMARTINV utilizes a fine-tuned LLaMa-7B model to infer invariants relevant to functional bug analysis and confirm bugs using validation algorithms. However, they are not consistently effective in identifying such bugs. Specifically, GPTScan cannot detect this issue as it fails during the GPT-based matching phase. SMARTINV generates invariants such as pricePerShare >  0 and balanceToShares(amount) >  0, which fail to capture the necessary details for analyzing this bug effectively.
PromFuzz introduces a novel approach that could effectively tackle and detect such bugs. Initially, PromFuzz employs an LLM-driven multi-perspective analysis that uses iterative questioning to pinpoint potential occurrences of this bug. Following this identification, PromFuzz extracts critical variables such as sender, allowances, and amount. An invariant checker is then constructed to verify the legitimacy of token quantity changes before and after transfers, which is integrated into the source code. Finally, PromFuzz involves bug-oriented fuzzing to ensure a comprehensive identification of functional bugs, completing the detection process.
1contract SimplePool{
2 // In this contract, we omit irrelevant code.
3 uint256 public pricePerShare;
4 function balanceToShares(uint256 balance) public view returns (uint256) {
5 return balance.mul(1e18).div(pricePerShare);
6 }
7 function transferFrom(address sender, address recipient, uint256 amount) public virtual override returns (bool) {
8 uint256 amountInShares = balanceToShares(amount);
9 _transfer(sender, recipient, amountInShares);
10 _approve(sender, _msgSender(), _allowances[sender][_msgSender()].sub(amountInShares,
11 "ERC20: transfer amount exceeds allowance"));
12 return true;
13 }
14}
Figure 1. The Approval Not Clear bug.
3. PromFuzz Design
In this section, we present the design details of PromFuzz. At a high level, PromFuzz aims to automatically find functional bugs in smart contracts. As shown in Figure 2, PromFuzz mainly consists of four modules. First, the LLM-driven multi-perspective analysis module accepts smart contracts as input and utilizes a dual-agent prompt engineering strategy, featuring the Auditor Agent and the Attacker Agent. These agents enable GPT to analyze smart contracts from distinct perspectives: a meticulous auditor’s viewpoint and a malicious attacker’s perspective, aiding in pinpointing potentially vulnerable functions accurately. Next, the invariant checker generation module leverages a dual-stage coupling approach, which focuses on generating invariant checkers that leverage logic information extracted from potentially vulnerable functions. Then, the bug-oriented analysis engine module strategically inserts the invariant checkers into specific potentially vulnerable functions, guiding the engine to explore these critical areas rather than searching aimlessly. Finally, the functional bug detection module executes comprehensive bug detection on smart contracts.
Figure 2. Framework of PromFuzz.
3.1. LLM-driven Multi-Perspective Analysis
The purpose of LLM-driven multi-perspective analysis is to analyze the high-level business logic concealed within smart contracts and pinpoint potentially vulnerable functions.
Detecting functional bugs in smart contracts requires a comprehensive understanding of the high-level business logic. Nevertheless, it is challenging to design a practical system that can automatically analyze the obscured business logic of smart contracts. The main problem lies in the fact that the high-level business logic is obscured within the low-level implementations of smart contracts, making it inaccessible through simple or hand-coded oracles that are commonly adopted by previous works.
Previous research has explored the use of LLMs to identify functional bugs in smart contracts by harnessing their significant emergent abilities. However, these studies often did not maximize the potential of LLMs, typically either inputting only high-level vulnerability descriptions or not fully engaging their capabilities. For example, Wang et al. (Wang et al., 2024) employed the LLaMA-7B model, but its relatively modest parameter size restricted its full emergent potential. Activating the emergent abilities of such models demands not only high-quality pre-training but also precise and effective instruction-following, which presents significant challenges. Furthermore, several previous works (Chen et al., 2023) (David et al., 2023) (Wang et al., 2024) have concentrated on enhancing the instruction-following capabilities of LLMs, frequently neglecting the crucial skill of step-by-step reasoning. This ability is vital for addressing complex problems, e.g., analyzing the obscured business logic of smart contracts, that require navigating through multiple layers of reasoning.
To address the above problem, we propose a novel LLM-driven analysis framework that employs a dual-agent prompt engineering strategy. This framework incorporates two specialized agents: the Auditor Agent and the Attacker Agent, each designed to enhance the emergent capabilities of GPT in different ways. Specifically, we regard the process of identifying bugs in smart contracts as an auditing task. Thus, the Auditor Agent is designed to act as a meticulous auditor, scanning smart contracts thoroughly and then detecting potential bugs. However, relying solely on the Auditor Agent might result in false positives and potentially overlook critical vulnerabilities, which could lead to irreversible damages if not identified early. To address this limitation, we introduce the Attacker Agent. This agent analyzes smart contract code from an attacker’s perspective, aiming to uncover exploitable weaknesses that the Auditor Agent might miss. This novel dual-agent architecture enables precise identification and reliable assessment of potential bugs, thereby enhancing the robustness of GPT’s output.
Both agents are equipped with precisely crafted prompts that aid in analyzing the obscured business logic of smart contracts. This dual-agent strategy ensures a thorough evaluation by encompassing both defensive and offensive aspects of security. By synthesizing the insights from both the Auditor Agent and the Attacker Agent, our framework significantly improves the precision of the analysis. It minimizes the removal of false positives while ensuring that critical bugs are accurately identified and retained for further action. In the following, we delve into the details of the dual-agent prompt engineering strategy employed in PromFuzz.
Figure 3. Prompt template for the Auditor Agent.
Auditor Agent Design. In the real world, smart contract auditors primarily concentrate on securing the core code and business operations of smart contracts. Inspired by this practical approach, we have designed our Auditor Agent to mirror the real-world auditing process. However, crafting the Auditor Agent is not trivial, as it involves a two-step process: initially capturing the essential code or business operations of smart contracts and subsequently identifying any abnormal behaviors within these core elements.
To address the above challenges, we design the Auditor Agent based on the CoT prompting strategy and the scenario and property matching methodology proposed by GPTScan (Sun et al., 2024). A “scenario” describes the code functionality under which a functional bug might occur, while a “property” details the vulnerable code attributes or operations. GPTScan initially selects ten representative functional bugs in smart contracts, following the classifications by Zhang et al. (Zhang et al., 2023). Upon further analysis, we observe that these categories could be streamlined for greater clarity and applicability.
Our refined approach consolidates the initial ten categories into four primary categories of functional bugs. Notably, we merge the categories of price manipulation by buying tokens and slippage into a single, more comprehensive category named non-AMM price oracle manipulation. Moreover, our manual analysis of the functional bugs leads to the identification of two new subcategories: improper handling of the deposit fee and wrong implementation of amount lock. We present the new classification results in Table 1.
Figure 3 outlines the prompt template for the Auditor Agent. First, GPT is informed that it is participating in a code auditing exercise. We then allow GPT to internally simulate the response formulation process five times and provide the most common answer to mitigate the randomness of the responses. Next, we input the target smart contract code and pose questions related to the functionality under which a functional bug might occur—i.e., the scenario. This question aids in identifying potential bug categories within the code. Finally, for any scenario where GPT confirms a potential issue, we follow up with specific questions about the vulnerable code attributes or operations—i.e., the property. If GPT responds affirmatively, we consider the presence of a functional bug; if not, we rule out that specific bug. The output of the Auditor Agent is fine-grained subcategories of bugs. Designing prompts requires only a one-time manual effort and can be easily extended to cover most functional bugs.
Figure 4. Prompt template for the Attacker Agent.
Attacker Agent Design. The Auditor Agent, although structured with a step-by-step reasoning process following conventional audit procedures, may lead to false positives and overlook critical bugs if solely relied upon. To address these limitations, we propose the Attacker Agent, which is specifically designed to emulate the cognitive processes of an attacker, actively searching for and exploiting bugs within smart contracts. By identifying and leveraging these weak points, the Attacker Agent enables GPT to conduct smart contract audits from an adversarial viewpoint. Designing the Attacker Agent poses significant challenges, as it requires a precise understanding and summarization of the manifestations and defensive measures associated with common logical vulnerabilities in smart contracts.
To address the above challenges, we gather an extensive dataset of real-world attack incidents for each bug category. Through detailed manual analysis of these incidents, we extract critical information, e.g., distinctive bug features, to craft effective prompts based on the CoT prompt strategy. Bug features refer to attributes or behaviors in the code that present security risks under certain conditions, which could potentially be exploited under specific conditions to carry out malicious activities. The process of utilizing the Attacker Agent involves several key steps, outlined in Figure 4. First, similar to the Auditor Agent, GPT is informed that it is participating in an ethical hacking exercise. Next, the target smart contract code is inputted, and GPT is questioned about potential bug features that could be exploited for malicious activities. Finally, for bugs whose features have been recognized during the last question, we provide the corresponding attack model for these bugs as prompts to GPT. We then query GPT whether the identified features could lead to exploitable attacks based on the given attack model. An affirmative response from GPT indicates the presence of a functional bug, while a negative response helps us dismiss that particular bug. The output of the Attacker Agent is coarse-grained primary categories of bugs.
Figure 5. Auditor-Attacker results fusion process corresponds to the motivating example in Figure 1.
Auditor-Attacker Results Fusion. To derive more accurate classifications of bugs, we integrate the results from both the Auditor Agent and the Attacker Agent. This fusion approach leverages the coarse-grained primary category insights from the Attacker Agent to refine and enhance the fine-grained subcategory determinations made by the Auditor Agent. The rules for results fusion are delineated as follows:
As shown in Algorithm 1, consider a scenario where for each analyzed code segment, the Attacker Agent identifies the primary bug categories $T$ at line 1, and the Auditor Agent identifies the subcategories $t$ at line 2. The fusion process operates under the following rules:
  1. 1)
    If $t$ is a subcategory of $T$, confirm $t$ as a candidate type (line 10-11).
  2. 2)
    In addition to confirming $t$, if $t$ belongs to $T$, also include all other subcategories of $T$ as candidate types to address potential overlaps or ambiguities (line 10-14).
  3. 3)
    If $t$ does not correspond to any subcategory of $T$, or if $T$ has not been identified, then exclude $t$ from the candidate types (line 17).
  4. 4)
    If $t$ is not determined but $T$ is identified, postpone the classification of this segment until further information is available or additional analysis is conducted (line 3).
For primary categories characterized by distinctly different subcategory features, such as insecure calculating logic and incorrect control mechanism, we apply rules 1), 3), and 4). For the other primary categories, we implement rules 2), 3), and 4). This selective application of fusion rules helps minimize false positives while maximizing the conversion of potential false negatives into true positives by the Auditor Agent. Although this method may initially increase the number of false positives, subsequent bug-oriented fuzzing processes are designed to effectively eliminate these inaccuracies.
Figure 5 illustrates an example of how our fusion rules are applied. First, the Auditor Agent identifies the code scenario as a token transfer and flags the property of this transfer as unsafe, classifying it under the subcategory of unauthorized transfer. Subsequently, the Attacker Agent analyzes potential bug features, indicating that these bugs could be exploited and categorizing them under the primary category of unauthorized behavior, which includes two subcategories: unauthorized transfer and approval not clear. Finally, we apply rule 2), 3) and 4) to integrate their analyses, leading to the identification of unauthorized transfer and approval not clear. The true positive identified in this code is approval not clear. Without this integration, we would fail to recognize this bug.
1 0:   $C$: smart contract code 0:   $B$: potential functional bug subcategory 1:  $T=Attacker\_Agent(C)$ 2:  $t=Auditor\_Agent(C)$ 3:  while $t\neq\varnothing$ do 4:     $t_{i}\leftarrow t.top()$ 5:     $t\leftarrow t-t_{i}$ 6:     while $T\neq\varnothing$ do 7:        $T_{j}\leftarrow T.top()$ 8:        $T\leftarrow T-T_{j}$ 9:        if $t_{i}$ is a subcategory of $T_{j}$ then 10:           $r_{mer}=Merge(T_{j},t_{i})$ 11:           $B=B.append(r_{mer})$ 12:           if  $j\in\{1,2\}$ then 13:              $r_{mut}=Mutation(T_{j},t_{i})$ 14:              $B=B.append(r_{mut})$ 15:           end if 16:        else 17:           Continue 18:        end if 19:     end while 20:  end while 21:  return  $B$
Algorithm 1 Auditor-Attacker Results Fusion
3.2. Invariant Checker Generation
While we have identified potentially vulnerable functions, we cannot solely rely on GPT’s results since its output includes randomness. To address this gap, it is crucial to develop a practical method that can minimize the impact of this randomness. One effective approach is to check the violation of invariants to further scrutinize the identified bugs. In smart contracts, an invariant is a condition or a set of conditions that always hold true regardless of the state of the contract at any point in its execution. These invariants are crucial for ensuring that the contract behaves correctly and securely, safeguarding against unintended actions and vulnerabilities. Therefore, there is a pressing need to implement a robust method for generating invariants to enhance functional bug detection. The primary difficulty in generating valid invariants lies in tailoring them specifically to different types of bugs. This requires a deep understanding of both the specific features of each bug and the context of the vulnerable code snippet.
Although SmartInv (Wang et al., 2024) utilizes a pre-trained LLaMA-7B model to infer invariants from smart contracts automatically, its generated invariants suffer from low-quality issues due to the following reasons. First, the model’s outputs frequently include hallucinations that significantly compromise the reliability of the generated invariants. Second, the presence of existing logical judgments in the code can obscure key logical information related to potential bugs, impacting the model’s ability to accurately generate relevant invariants. Third, the small parameter size of the LLaMA-7B model necessitates a well-designed dataset for effective fine-tuning to improve its performance in generating precise and useful invariants.
To address these problems, we design a dual-stage coupling method to generate invariant checkers that take logic information extracted from potentially vulnerable functions as input and generate corresponding invariants. Specifically, in the first stage, we implement a hierarchical matching approach that utilizes GPT’s capabilities to extract essential variables and statements. This preliminary step focuses on gathering critical information rather than generating complete invariants directly. In the second stage, we introduce a template-based checker generation approach. We design $6$ invariant checker templates specifically tailored to integrate critical variables and principal statements extracted in the first stage. This coupling ensures the generated invariants are both relevant and robust. In the following, we present the details of the invariant checker generation.
Table 2. Characteristics of critical variables and principal statements for each bug subcategory, along with the associated invariant checkers.
Bug Subcategory Key Variables (V) & Statements (S) Invariant Checker
AMM Price Oracle Manipulation V1: calculated price of the LP token PriceChange_Checker(V1)
Non-AMM Price Oracle Manipulation V1: tokens spent by the user during the swap process ExchangeRate_Checker(V1,V2)
V2: the amount of tokens received by the user after the swap
Approval Not Clear V1: sending address for token transfers TokenChange_Checker(V1,V2,V3)
V2: authorization array for token transfers
V3: the amount of tokens being transferred
Unauthorized Transfer V1: sending address for token transfers TokenChange_Checker(V1,V2,V3)
V2: authorization array for token transfers
V3: the amount of tokens being transferred
Wrong Checkpoint Order S1: the first to call Checkpoint StatementOrder_Checker(S1,S2)
S2: the first to perform a calculation or update on the tokens
Risky First Deposit V1: the total supply of tokens ShareSafety_Checker(V1,V2,V3)
V2: the share that a user can obtain
V3: tokens spent by the user
Wrong Interest Rate Order S1: the first to update the interest rate StatementOrder_Checker(S1,S2)
S2: the first to use the interest rate to calculate or update the
token balance, rewards, or shares
Improper Handling of Deposit Fee V1: tokens spent by the user without any deductions TokenChange_Checker(V1,V2,V3)
V2: the net amount added to the user’s balance after deducting
the deposit fee
V3: the amount of deposit fee
Wrong Implementation of Amount Lock
V1: the array where the first element is used to store the user’s
original locked amount
TokenChange_Checker(V1,V2,V3)
V2: the amount by which the user’s locked funds have increased
or decreased
V3: stores the new locked amount after increasing or decreasing
based on the isIncrease flag
Vote Manipulation V1: contract variable for voting tokens StateChange_Checker(V1,V2,S1)
V2: the number of votes cast by the user in this round
S1: the first to execute a voting action
Critical Variables and Principal Statements Extraction. To develop the invariant checker, it is essential to initially identify and extract critical variables and principal statements related to the business logic from smart contracts. These elements are important in determining the invariants associated with different types of bugs, which are then used to populate a predefined checker template. However, the identification of these critical elements poses significant challenges. It requires a thorough understanding of the invariants pertinent to various types of bugs and the design of a practical and effective method for their extraction from smart contracts.
In response to this challenge, we propose a hierarchical matching method that capitalizes on the capabilities of GPT for the extraction of critical variables and principal statements. Our approach unfolds in two primary steps. First, we perform an in-depth analysis of each bug category to pinpoint critical variables and principal statements that are vital for comprehending their logic. For instance, in the scenario of an AMM price oracle manipulation bug, a critical variable is the one that holds the calculated price of the LP token. We present the characteristics of critical variables and principal statements for each bug subcategory in Table 2. Second, leveraging the identified elements, we formulate specific prompts and input them into GPT to query each vulnerable function as identified by the LLM-driven multi-perspective analysis module. Figure 6 illustrates the prompt template designed to extract critical variables and principal statements from code snippets potentially vulnerable to AMM price oracle manipulation. For different bug categories, we can substitute [%Critical Variable%] and [%Principal Statement%] in the template with the predefined elements specific to those bugs. With these tailored prompts, GPT is able to precisely pinpoint the relevant critical variables and principal statements, enhancing the accuracy of our bug detection process.
Figure 6. Prompt template for extracting critical variables from code snippets potentially vulnerable to AMM price oracle manipulation.
Template-based Checker Generation. While we have extracted critical variables and principal statements from smart contracts, we currently lack an efficient method to convert these into usable invariants for bug detection. To fill this gap, we propose a template-based checker generation method. Specifically, for each type of bug, we design a corresponding invariant checker template that targets the vulnerable features of different bugs, as detailed in Table 2. We mainly design 6 types of checkers: PriceChange_Checker, ExchangeRate_Checker, TokenChange_Checker, StatementOrder_Checker, ShareSafety_Checker, and StateChange_Checker. Each checker includes a conditional statement that assesses the contract state. For instance, as shown in Figure 7, the PriceChange_Checker for AMM price oracle manipulation determines whether the variable (i.e., calculated price of the LP token) falls below $90\%$ or exceeds $110\%$ of the old price. Such a condition indicates that the price change surpasses the ±10% threshold, typically signaling abnormal price fluctuations and triggering an alert for an existing AMM price oracle manipulation bug. These checkers could be dynamically combined to comprehensively cover a broad range of functional bugs, ensuring scalability.
Figure 7. The checker for AMM price oracle manipulation.
3.3. Bug-Oriented Analysis Engine
Though we have identified potentially vulnerable functions in smart contracts and generated the corresponding invariant checkers, we still lack a practical method to leverage them for functional bug detection. Existing works mainly rely on static analysis tools to validate rule violations or invariant breaches to detect functional bugs. For instance, GPTScan leveraged Falcon (met, 2024), a static analysis tool built upon Slither (Feist et al., 2019), to confirm the existence of the candidate vulnerable functions. Similarly, SMARTINV employed a customized static checker based on VERISOL (Wang et al., 2019a) to detect violations of inferred invariants. Nevertheless, these tools often suffer from false positives and cannot fully simulate contract behavior in real-world environments, potentially missing runtime-specific issues, leading to false negatives. To tackle the above problem, we implement a bug-oriented analysis engine with the following designs.
Invariant Checker Insertion. Inserting invariant checkers into smart contracts requires careful consideration to avoid false negatives caused by incorrect or inappropriate insertion points. We propose two criteria for strategic insertion: the variable-oriented invariant insertion criterion and the statement-oriented invariant insertion criterion. The variable-oriented criterion focuses on inserting invariants before and after critical code blocks that manipulate critical variables, ensuring that the operations conform to the invariants. Similarly, the statement-oriented criterion places invariants around crucial statements to verify if their execution logic and sequence adhere to the defined invariants. Utilizing these criteria, invariant checkers are automatically positioned in optimal locations within the code, taking into account the logical relationships between critical variables, principal statements, and their positions in the code block.
Bug-Oriented Fuzzing. After inserting invariant checkers into smart contracts, we then employ fuzzing to address the aforementioned limitations in static analysis methods. Traditional dynamic analysis techniques often explore programs randomly without a specific target, typically identifying bugs through program crashes. However, functional bugs in smart contracts do not generally lead to crashes and are less likely to be detected through random testing. To mitigate this, we implement a bug-oriented fuzzing approach that directs the fuzzer to specifically target functions with inserted invariant checkers, rather than performing random exploration. Our method builds upon ItyFuzz (Shou et al., 2023), a state-of-the-art hybrid fuzzer for smart contracts that integrates symbolic execution with fuzzing. ItyFuzz has shown great performance in finding implementation bugs but falls short in detecting functional bugs based on our evaluation in Section 4.4. Therefore, to bridge the gap, we customize ItyFuzz by leveraging its functionality that allows verification of custom invariants. This customization involves inserting a specific bug alert function into each invariant checker, which allows us to direct the fuzzer to activate these targeted checkers rather than engaging in random exploration. Our method significantly improves the efficiency and effectiveness of the fuzzing process.
0:   $C$: smart contract code $B$: potential functional bug subcategory 0:   $R$: the set of detected functional bug $Q$: the set of results requiring further inspection 1:  for each $B_{i}\in B$ do 2:     $v,s=Extract\_Variables\_and\_Statements(B_{i},C)$ 3:     $inv=Generating\_invariants(v,s)$ 4:     $CwI=Invariant\_Checker\_Insertion(inv,B_{i},C)$ 5:     $C_{s}\leftarrow initial\_infant\_state\_corpus$ 6:     $C_{ts}\leftarrow initial\_transaction\_and\_state\_pair\_corpus$ 7:     while $t\neq\varnothing$ do 8:        $s_{mut},t_{mut}\leftarrow input\_mutation(C_{ts},C_{s})$ 9:        $f,s^{\prime},violation\leftarrow Execution(s_{mut},t_{mut},CwI)$ 10:        if  $violation\ is\ found$  then 11:           $R=R.append(inv,B_{i})$ 12:        end if 13:        $update\_corpus(f,s^{\prime},s_{mut},t_{mut})$ 14:     end while 15:     if  $violation\ is\ not\ found$  then 16:        $Q=Q.append(inv,B_{i})$ 17:     end if 118:  end for 19:  return  $R$, $Q$
Algorithm 2 Functional Bug Detection
3.4. Functional Bug Detection
The functional bug detection module aims to perform comprehensive bug detection on smart contracts. Algorithm 2 outlines the procedure for applying this methodology. Specifically, in this module, we first compile the target smart contracts with strategically inserted invariant checkers. Next, we perform fuzzing on these smart contracts. When the fuzzer activates an invariant checker and detects a discrepancy, it signals the presence of a bug. According to the results, we then generate a comprehensive bug report that documents all identified functional bugs within the smart contracts. This report serves as a crucial tool for developers to understand and rectify bugs in their smart contracts.
4. System Evaluation
In this section, we evaluate the performance and capabilities of PromFuzz from multiple perspectives. Our evaluation is structured around the following key research questions:
RQ1: How does the dual-agent prompt engineering strategy contribute to the performance of PromFuzz? (Section 4.2)
RQ2: What is PromFuzz’s accuracy in generating invariant checkers? (Section 4.3)
RQ3: How effective is PromFuzz at detecting functional bugs in smart contracts? (Section 4.4)
RQ4: Is PromFuzz capable of discovering zero-day functional bugs in real-world smart contracts? (Section 4.5)
4.1. Dataset
To ensure a comprehensive assessment of PromFuzz’s effectiveness, we create the following 6 datasets covering a diverse range of functional bug categories. These datasets are sourced from DeFiHackLabs (def, 2024), Web3bugs (Zhang et al., 2023), GPTScan (Sun et al., 2024), on-chain smart contracts and real-world DeFi projects. We selectively use these datasets, excluding projects lacking necessary external files or configuration information for compilation, as PromFuzz’s dynamic analysis necessitates complete setups. Furthermore, datasets like SmartBugs (Durieux et al., 2020) and VeriSmart (So et al., 2020), which lack functional bugs, are excluded from our evaluation.
Bug-free smart contract dataset ($D_{BF-SC}$) includes well-audited and widely used smart contracts that are free from any known functional bugs. These contracts are sourced from GPTScan (Sun et al., 2024) and serve as a benchmark to assess the false positive rate of PromFuzz. Specifically, $D_{BF-SC}$ consists of $303$ open-source contract projects from $6$ mainstream Ethereum-compatible chains.
Exploited smart contract dataset ($D_{Exploit-SC}$) includes smart contracts that have been subjected to real-world attacks in the past, each causing significant monetary losses. $D_{Exploit-SC}$ is sourced from DeFiHackLabs (def, 2024) and contains $7$ real-world attack cases that we have manually verified as being caused by functional bugs. Among these, $4$ cases are due to price oracle manipulation, $2$ are due to unauthorized behavior, and $1$ is due to incorrect control mechanism.
Bug-verified smart contract dataset ($D_{BV-SC}$) includes smart contracts audited through platforms like Code4rena (Cod, 2024) and Immunefi (Imm, 2024), with confirmed high-risk functional bugs by the developers. $D_{BV-SC}$ is sourced from Web3bugs (Zhang et al., 2023), which includes $10$ functional bugs across $36$ smart contract files. Among these, $5$ bugs result from insecure calculating logic, $2$ are due to incorrect control mechanisms, $2$ are due to unauthorized behavior, and $1$ is due to price oracle manipulation.
Synthetic bug-injected smart contract dataset ($D_{Synth-SC}$) includes smart contracts with manually injected functional bugs. The purpose of this dataset is to enlarge $D_{Exploit-SC}$ and $D_{BV-SC}$, providing a comprehensive evaluation of PromFuzz. We create $D_{Synth-SC}$ by manually injecting functional bugs into well-performing, bug-free projects on the Ethereum-compatible chain, which includes $6$ manually injected functional bugs across $44$ smart contract files. Among these, $4$ bugs are due to unauthorized behavior and $2$ are due to incorrect control mechanisms.
Symbol-extracted smart contract dataset ($D_{SE-SC}$) includes smart contact functions with manually extracted and verified critical variables and principal statements. Specifically, $D_{Extracted-SC}$ includes $55$ extracted critical variables and principal statements across $23$ smart contract functions.
Real-world DeFi project dataset ($D_{Real-DeFi}$) comprises DeFi projects currently undergoing public audits through various platforms. Specifically, $D_{Real-DeFi}$ includes $6$ such DeFi projects.
Table 3. False positive rate evaluation.
Tool $D_{BF-SC}$
FPR (Function-Level) FPR (File-Level)
PromFuzz 6.31% 15.38%
GPTScan 7.16% 15.75%
SMARTINV N/A 72%
Table 4. Comparison of Dual-agent and Auditor-agent.
Tool $D_{Exploit-SC}$ $D_{BV-SC}$ $D_{Synth-SC}$ Overall
TP FP FN TP FP FN TP FP FN TP FP FN Recall F1-Score
Dual-agent 6 16 1 9 12 1 6 6 0 21 34 2 91.30% 53.85%
Auditor-agent 5 9 2 7 10 3 3 5 3 15 24 8 65.22% 48.39%
4.2. Dual-Agent Architecture Evaluation
This subsection addresses RQ1 by assessing the effectiveness of the dual-agent architecture of PromFuzz. We begin with evaluating the false positive rate of the LLM-driven multi-perspective module in PromFuzz. While PromFuzz generally mitigates false positives during the fuzzing phase, the occurrence of additional false positives requires PromFuzz to expend more resources on verifying non-existent bugs. These unnecessary validation operations will incur additional costs in generating invariant checkers and consume substantial time during the fuzzing phase. To this end, we compare PromFuzz with two established tools: GPTScan and SMARTINV. GPTScan integrates GPT with static analysis to detect functional bugs in smart contracts. SMARTINV leverages a fine-tuned LLM to infer invariants from smart contracts and performs further analysis based on a static tool. In this experiment, PromFuzz is exclusively compared with the LLM modules of GPTScan and SMARTINV. We set up both tools according to the methodologies described in their respective papers and then perform them on $D_{BF-SC}$. Both GPTScan and PromFuzz utilize GPT-4-turbo as their foundational model.
Next, we conduct an ablation study to evaluate the performance differences between the dual-agent based PromFuzz and a version that utilizes only the Auditor Agent. To achieve this goal, we construct two new PromFuzz, the first one is PromFuzz + Dual-agent, and the second one is PromFuzz + Auditor-agent. This comparison is conducted using three datasets: $D_{Exploit-SC}$, $D_{BV-SC}$, and $D_{Synth-SC}$.
False Positive Rate Evaluation. Both PromFuzz and GPTScan can analyze whether a function contains functional bugs, operating at the function-level, while SMARTINV is limited to file-level analysis. For a fair comparison, we conduct evaluations at both the function and file levels. As illustrated in Table 3, PromFuzz has a slight advantage with false positive rates of $6.31\%$ and $15.38\%$, respectively, than GPTScan. Despite strictly adhering to the guidelines for SMARTINV and utilizing its publicly released code, SMARTINV records a significantly higher false positive rate of $72\%$ at file-level.
We further analyze the false positives of these tools. For SMARTINV, the primary focus of its LLaMA-7B fine-tuning strategy and prompt design is on generating critical invariants. However, this approach lacks comprehensive reasoning capabilities necessary for effective bug detection. This deficiency in the analytical strategy significantly contributes to the tool’s notably high false positive rate. For GPTScan, the primary reason for false positives arises from its single-agent design, which can lead to the incorrect identification of correct behaviors as functional bugs, such as those involving insecure calculating logic. Although PromFuzz has similar reasons, its innovative dual-agent design effectively mitigates a greater number of false positives compared to single-agent systems.
Ablation Study. As shown in Table 4, Dual-agent shows substantial performance improvements over the single-agent version. Specifically, on the $D_{Exploit-SC}$ dataset, Dual-agent identifies 6 true positives, 16 false positives, and only 1 false negative. On the $D_{BV-SC}$ dataset, it detects 9 true positives, 12 false positives, and 1 false negative. Notably, on the $D_{Synth-SC}$ dataset, Dual-agent identifies 6 true positives, 6 false positives, and no false negatives. Overall, Dual-agent achieves a recall of 91.30% and an F1-score of 53.85%, identifying 21 true positives and 34 false positives across all datasets, with only 2 false negatives. In contrast, Auditor-agent detects fewer true positives and has more false negatives, achieving a recall of 65.22% and an F1-score of 48.39%.
By adopting our dual-agent prompt engineering strategy, PromFuzz + Dual-agent significantly improves true positives and eliminates false negatives. Although PromFuzz + Dual-agent exhibits a higher number of false positives, these are manageable as they can be effectively filtered out during the subsequent fuzzing phase.
We further analyze the false positives and false negatives by PromFuzz + Dual-agent and PromFuzz + Auditor-agent. For PromFuzz + Dual-agent, though the Attacker Agent increases true positives by considering multiple angles, it still has limitations and can lead to over-generation of false positives. For PromFuzz + Auditor-agent, auditor’s perspective, which focuses solely from an auditing standpoint, can result in underreporting, filtering out some high-risk vulnerabilities. The limitations of auditor prompt rules, which, like those of the attacker, cannot cover all bugs, resulting in some being missed. Besides, for both PromFuzz + Dual-agent and PromFuzz + Auditor-agent, GPT’s understanding is still insufficient, leading to cases of prompt misinterpretation.
4.3. Invariant Checker Generation Accuracy
This subsection answers RQ2. In this step, we mainly evaluate the accuracy of PromFuzz in extracting critical variables and principal statements from potentially vulnerable functions, which determines the accuracy of final invariant checkers. We deploy PromFuzz on dataset $D_{SE-SC}$ for this evaluation.
As shown in Table 5, the invariant checker generation module demonstrates excellent performance in critical variable and principal statement extraction, achieving an overall precision of $98.18\%$. Specifically, for the variable extraction task, PromFuzz attains a perfect precision of $100\%$. For the statement extraction task, it achieves a precision of $83.33\%$, with only one false positive. This false positive arises from the complex business behaviors in smart contracts, where PromFuzz fails to differentiate between multiple similar behaviors in a single complex case. These critical variables and principal statements are expected to derive $23$ correct invariant checkers but ultimately generate $22$ correct ones, resulting in a precision of $95.65\%$.
Table 5. Variables and statements extraction accuracy.
Task $D_{SE-SC}$
TP FP Precision
Variables Extraction 49 0 100%
Statements Extraction 5 1 83.3%
Overall 54 1 98.18%
Table 6. Functional bug detection accuracy.
Tool $D_{Exploit-SC}$ $D_{BV-SC}$ $D_{Synth-SC}$ Overall
TP FP FN TP FP FN TP FP FN TP FP FN Recall F1-Score
PromFuzz 6 0 1 8 0 2 6 0 0 20 0 3 86.96% 93.02%
GPTScan 2 4 5 5 4 5 0 4 6 7 12 16 30.43% 33.33%
SMARTINV 4 19 3 1 20 9 1 4 5 6 43 17 26.09% 16.67%
ItyFuzz 0 0 7 0 0 10 0 0 6 0 0 23 0 0
SMARTIAN 0 0 6 1 0 8 2 0 2 3 0 16 15.79% 27.27%
4.4. Functional Bug Detection Accuracy
This subsection answers RQ3. In this step, we evaluate the functional bug detection accuracy of PromFuzz with two metrics: recall and F1-score. We compare PromFuzz with four off-the-shelf tools: GPTScan, SMARTINV, ItyFuzz, and SMARTIAN (Choi et al., 2021). We do not compare PromFuzz with several program verification-based techniques (Kalra et al., 2018; Permenev et al., 2020; Duan et al., 2022), since they are largely ineffective in detecting functional bugs, as their constraints are designed for implementation bugs. Both ItyFuzz and SMARTIAN are fuzzers designed for detecting bugs in smart contracts. We perform them on $D_{Exploit-SC}$, $D_{BV-SC}$, and $D_{Synth-SC}$, respectively.
As shown in Table 6, PromFuzz demonstrates superior performance in detecting functional bugs across three datasets. On $D_{Exploit-SC}$, PromFuzz identifies 6 true positives with no false positives and only 1 false negative. Similarly, on $D_{BV-SC}$, it identifies 8 true positives, with no false positives and 2 false negatives. For $D_{Synth-SC}$, PromFuzz perfectly identifies 6 true positives with no false negatives or false positives. Cumulatively, PromFuzz achieves an overall recall of 86.96% and an F1-score of 93.02%, detecting 20 true positives and maintaining 0 false positives across all datasets, with only 3 false negatives. Further analysis of the false negatives identified by PromFuzz reveals two primary causes. First, one false negative results from inaccuracies in the extraction of critical variables and principal statements. Second, the remaining two false negatives occur during the LLM-driven analysis step, where GPT fails to identify two critical bugs. Besides, with each smart contract project scan, PromFuzz takes an average of 43.8 seconds and costs approximately \$0.15, demonstrating its great efficiency and cost-effectiveness.
In comparison, GPTScan and SMARTINV yield similar numbers of true positives and false negatives. However, SMARTINV records the highest number of false positives among the five tools assessed. Furthermore, ItyFuzz fails to identify functional bugs without the specific invariants we provide. SMARTIAN, another fuzzing tool, shows inconsistent performance as it fails to operate on several smart contract files. Consequently, we exclude SMARTIAN from testing against all bugs. We further analyze the false positives and false negatives of these tools. The reasons behind the false positives for both GPTScan and SMARTINV are already discussed in Section 4.2. The false negatives for GPTScan mainly arise because its prompts do not fully cover several specific functional bugs, such as unauthorized behavior, despite considerable efforts. SMARTINV experiences false negatives due to its lack of comprehensive reasoning capabilities, resulting in some bugs being overlooked. Both ItyFuzz and SMARTIAN are limited by their original oracles, which lack the capability to identify functional bugs effectively.
Table 7. Zero-day functional bug detection results. TVL represents the total value of cryptocurrency assets locked in the respective DeFi projects.
DeFi Project Type TVL # Zero-day # CVE
Project Bug
1 Yield & Lengding \$ 136.8M 7 7
2 Yield \$ 14,300M 3 3
3 Yield \$ 0.172M 4 4
4 Exchanges \$ 102M 5 5
5 Yield & Sericves \$ 3,700M 5 5
6 Exchanges \$ 1.79M 6 0
Total \$ 18.2B 30 24
4.5. Zero-day Functional Bug Detection
This subsection answers RQ4. We perform PromFuzz on $D_{Real-DeFi}$ to evaluate its performance in detecting zero-day functional bugs in real-world smart contracts. As illustrated in Table 7, PromFuzz successfully identifies $30$ zero-day functional bugs across $6$ DeFi projects. We have reported all detected bugs to the respective vendors, and $24$ of these have been assigned CVE IDs. The total market value of DeFi projects affected by these bugs is estimated at \$$\$18.2$ billion. As these bugs have not yet been fully resolved, we maintain confidentiality by anonymizing the details of both the vendors and the specific CVE information. Moreover, we notice that zero-day bugs primarily fall into four categories: 13 bugs involve AMM price oracle manipulation, $7$ concern non-AMM price oracle manipulation, $4$ are related to wrong checkpoint order, $4$ pertain to risky first deposit, and $2$ involve wrong interest rate order. Further analysis in Section 5 delves into the practical impacts of these bugs, examining specific cases to illustrate how attackers exploit these bugs.
5. Case Study
To obtain an in-depth understanding of the practical impacts of functional bugs, we detail two such bugs identified by PromFuzz in DeFi projects.
1contract SimplePool{
2 function _deposit(address _recipient, uint256 _amount) internal returns (uint256) {
3 totalValueInLp += _amount;
4 uint256 share = _sharesForDepositAmount(_amount);
5 if (_amount > type(uint128).max || _amount == 0 || share == 0){
6 revert InvalidAmount();
7 }
8 eETH.mintShares(_recipient, share);
9 return share;
10 }
11 function _sharesForDepositAmount(uint256 _depositAmount) internal view returns (uint256) {
12 uint256 totalPooledEther = getTotalPooledEther() - _depositAmount;
13 if (totalPooledEther == 0) {
14 return _depositAmount;
15 }
16 return (_depositAmount * eETH.totalShares()) / totalPooledEther;
17 }
18}
Figure 8. The Risky First Deposit bug (line 13-15).
Risky First Deposit. As shown in Figure 8, the code contains a risky first deposit bug that can lead to an unfair advantage for the first user and a distorted distribution of shares. Specifically, at line 14, when totalPooledEther == 0, the share variable at line 5 is set equal to _depositAmount, since the return value is assigned to _depositAmount in the _sharesForDepositAmount() function. Subsequent depositors have their shares calculated by (_depositAmount $\times$ eETH.totalShares()) / totalPooledEther, which depends on the values of eETH.totalShares() and totalPooledEther. While the calculation logic seems standard for compliant depositors, the first depositor can exploit the system by manipulating totalPooledEther. We can front-run other depositors’ transactions and inflate the price of pool tokens through a substantial “donation.” For instance, a malicious early user could deposit 1 wei of asset tokens as the first depositor, receiving 1 wei of shares tokens. Subsequently, the malicious sends 1 ETH to the pool, inflating totalPooledEther to 1 ETH + 1 wei (equal to ($11e18+$ 1) wei). In this scenario, the second depositor, depositing 2 ETH (equal to $182e$ 18 wei) of asset tokens, would receive only 1 wei of shares tokens. Because 1.99 wei of shares tokens (calculated by $2e18/(1e18+1)$) will be rounded down to 1 wei of shares tokens in smart contracts. They will lose 0.5 ETH if they withdraw right after the _deposit(). This disproportionate calculation of the first user’s shares can disadvantage subsequent users who receive fewer shares relative to their deposits.
AMM Price Oracle Manipulation. As shown in Figure 9, the code contains an AMM price oracle manipulation bug, which arises from the method of calculating token prices based solely on the supplies of two tokens. Specifically, _reserveA and _reserveB represent the respective reserve amounts of two assets. In line 15 of the code, the amount of token B is calculated using the amount of token A, along with the reserves of token A and token B. However, if the reserve amount of either token is maliciously manipulated, it can distort the price and lead to significant price fluctuations. For instance, consider token exchanges in the smart contract using the PoolHelpers library. Initially, the pool contains 500 token A and 500 token B. An attacker exchanges 100 token A for 100 token B, changing the pool to 600 token A and 400 token B. Then, the second user exchanges 30 token A and can only get 20 token B in return, leaving the pool with 630 token A and 380 token B. Finally, the attacker uses the 100 token B they obtained to exchange for 165 token A, ultimately making a profit of 65 token A.
1library PoolHelpers {
2 function quote(uint256 _amountA, uint256 _reserveA, uint256 _reserveB) internal pure returns (uint256 amountB_) {
3 require(_amountA != 0, "SmardexHelper: INSUFFICIENT_AMOUNT");
4 require(_reserveA != 0 && _reserveB != 0, "SmardexHelper: INSUFFICIENT_LIQUIDITY");
5 amountB_ = (_amountA * _reserveB) / _reserveA;
6 }
7}
8contract SimplePool{
9 address tokenA,tokenB;
10 function swap(uint256 _amountA,address _tokenA) return (uint256 _amountB){
11 require(_tokenA==tokenA||_tokenA==tokenB);
12 address _tokenB = (_tokenA==tokenA)?tokenB:tokenA;
13 _reserveA = getReservers(_tokenA);
14 _reserveB = getReservers(_tokenB);
15 _amountB = PoolHelpers.quote(_amountA, _reserveA, _reserveB);
16 }
17}
Figure 9. The AMM Price Oracle Manipulation bug (line 15).
6. Discussion
While PromFuzz achieves significant results in detecting functional bugs in smart contracts, it has several limitations. First, some false negatives occur due to inaccuracies in extracting critical variables and the inherent randomness in the LLM-driven analysis. Despite these issues, PromFuzz still outperforms existing methods in uncovering deeper functional bugs. Additionally, our bug-oriented analysis currently targets four main categories with ten subcategories, potentially missing some complex bug types. In the future, we plan to fine-tune a specific LLM for enhanced detection of functional bugs in smart contracts and develop new prompts to cover a wider range of bug categories, thus improving PromFuzz’s comprehensiveness. Moreover, since our analysis partially relies on ItyFuzz, which is not specifically designed for functional bugs and lacks full utilization of contract status information, future efforts will focus on better integrating LLMs with dynamic analysis techniques to more effectively and comprehensively utilize contract status information, aiming to capture deeper functional bugs.
7. Related work
Recent efforts to evaluate the security of smart contracts have produced numerous methodologies. These approaches are broadly classified into three categories: static analysis, dynamic analysis, and AI-based analysis.
Static Analysis. Tools like Slither (Feist et al., 2019), Zeus (Kalra et al., 2018), Securify (Tsankov et al., 2018), Madmax (Grech et al., 2018), Invcon (Liu and Li, 2022), Verismart (So et al., 2020), and Verisol (Wang et al., 2019a) analyze smart contract code statically. Slither converts Solidity code into SlithIR to find bugs. Securify uses compliance and violation patterns to assess code safety. Madmax decompiles EVM bytecode and uses a logic-based specification for bug detection. Verisol translates Solidity into Boogie, formalizing semantic conformance against state machine workflows and reducing semantic checking to safety verification. However, these tools primarily focus on implementation bugs like integer overflow and reentrancy, and are less effective at identifying functional bugs.
Dynamic Analysis. Tools like ItyFuzz (Shou et al., 2023), Smartian (Choi et al., 2021), Harvey (Wüstholz and Christakis, 2020), Echidna (Grieco et al., 2020), ContractFuzzer (Jiang et al., 2018), Mythril (myt, 2024), Sailfish (Bose et al., 2022), Manticore (Mossberg et al., 2019), Teether (Krupp and Rossow, 2018), and Oyente (Luu et al., 2016) analyzes programs during execution. ItyFuzz uses snapshot-based fuzzing to reduce re-execution overhead, prioritizing intriguing snapshot states. Smartian combines static analysis for seed generation with feedback mechanisms. Sailfish identifies state-inconsistency bugs using a combination of lightweight exploration and symbolic evaluation. While dynamic analysis captures more realistic behaviors by executing smart contracts, it still primarily targets implementation bugs and struggles to effectively detect functional bugs.
AI-based Analysis. SmarTest (So et al., 2021), ILF (He et al., 2019),xFuzz (Xue et al., 2024), and GNN-based tools (Zhuang et al., 2020; Liu et al., 2021, 2023) use AI to enhance smart contract bug detection. SmarTest combines symbolic execution with a language model to prioritize bugs. GNN-based tools leverage graph neural networks with expert patterns. While these tools improve detection capabilities, they mainly address implementation bugs. Recently, LLM-based tools like GPTScan (Sun et al., 2024) and SMARTINV (Wang et al., 2024) have been used to address functional bugs by analyzing discrepancies between business logic and code. These works, including the PromFuzz introduced here, open new avenues for detecting functional bugs in smart contracts using AI’s cognitive capabilities.
8. Conclusion
In this paper, we introduce PromFuzz, an automated and practical system designed to detect functional bugs in smart contracts, which bridges the gap between understanding the high-level business model logic and scrutinizing the low-level implementations in smart contracts. PromFuzz achieves $86.96\%$ recall and $93.02\%$ F1-score in detecting functional bugs, representing a significant performance improvement over state-of-the-art methods. Additionally, we have created the first ground-truth dataset for functional bugs, including $23$ verified functional bugs from various DeFi projects. Furthermore, our in-depth analysis of $6$ real-world DeFi projects has identified $30$ zero-day bugs, with $24$ of them receiving CVE IDs. Our discoveries have safeguarded assets totaling \$$\$18.2$ billion from potential monetary losses.